OpenStudy (anonymous):

use the point-slope form of the equation to indentify a piont the passes through and the slope of the line: y+4=-2(x+1)

OpenStudy (anonymous):

y-y1=m(x-x1) ---> point slope formula point that passes through is (x1,y1) and slope is m therefore your point would be (-1,-4) and the slope m=-2 if you plug this into the general point slope formula you get: y-(-4)=-2(x-(-1)) which is the same thing as y+4=-2(x+1)